# Reward-Based Collision-Free Algorithm for Trajectory Planning of Autonomous Robots

### v1

@@ IV-B1 Ground Vehicle
The maximum available reward is given by $\sum_{i=2}^{11}w_{\boldsymbol{q}_{i}}=110$. Due to the distance constraint, only a subset of the waypoints can be reached, resulting in a maximum collected reward of $92$. The result is illustrated by the trajectory shown in Fig. 14b. Hypothetically, instead of squeezing itself through two obstacles to reach the waypoint with reward 8, it reaches the waypoints of reward 14, 12, and 18, keeping the other waypoints in the same order. In this case, the total collected reward would increase to 96; but the distance traveled would be 8.5 m $>d_{\mathrm{max}}$.
@@ IV-B2 Quadrotor
The geometric environment and obstacles are shown in Fig. 15. There is no distance constraint but a mission time window $t_{\mathrm{max}}=30$ s. The speed constraint is $|\dot{\boldsymbol{x}}|\leq 0.3$ m/s, and the maximum linear acceleration is $0.5$ m/s2. In total, 11 intermediate waypoints are distributed, with rewards starting at $w_{\boldsymbol{q}_{2}}=2$ and increasing in increments of 2, culminating in $w_{\boldsymbol{q}_{12}}=22$. Only the initial position is specified, allowing the final point to remain free. The optimal trajectory, shown in Fig. 15, achieves a total reward of $96$ within a time of $29.8$ s.
@@ IV-B3 Quadruped
We demonstrate our approach using the Unitree Go1 quadruped robot. The robot is equipped with two front and two rear legs, each having three joints with specific angular constraints. To the best of the authors’ knowledge, no differentially flat mapping from the states and inputs of the robot’s contact-aware multi-rigid-body dynamic model to its outputs has been presented in the literature. Since the proposed algorithm emphasizes high-level task allocation and planning, we utilize a simplified kinematic model that takes forward-backward speed, left-right speed, and orientation as inputs, following a similar concept in the literature [39]. Although this simplification limits more complex planning, such as navigating uneven terrain, the use of hierarchical planning with models of varying fidelity remains effective for task planning [39]. In this high-level framework, we consider the state $\boldsymbol{x}=[x_{1}\ x_{2}\ \theta]^{T}$ in a global reference frame, as indicated in Fig. 16, and we define the flat output as $\boldsymbol{y}=\boldsymbol{x}$. Then, the map from the flat output to the state is the identity, and the map from flat output and its derivatives to the flat inputs is
@@ V-A1 Permutation-Based Integer Formulation
for $i=2,3,\dots,\kappa-1$ (waypoints), and $j=2,3,\dots,\kappa-1$ (positions in the sequence). We exclude the fixed initial and final waypoints. In order to have each waypoint assigned to at most one position, one can use $\sum_{j=2}^{\kappa-1}z_{ij}\leq 1,\forall i\in\{2,3,\dots,\kappa-1\}$, and similarly for each position to have at most one waypoint assigned $\sum_{i=2}^{\kappa-1}z_{ij}\leq 1,\forall j\in\{2,3,\dots,\kappa-1\}$. With these two constraints we assure there is no repetitions, and the robot will reach single waypoint at a time. Moreover, since we want to force variable length sequences, we include the total number of assignments (intermediate waypoints) as a decision variable $\lambda$, and we constraint the waypoint assignments by $\sum_{i=2}^{\kappa-1}\sum_{j=2}^{\kappa-1}z_{ij}-\lambda\leq 0$. Finally, defining the decision variable $\mathbf{x}\in\mathbb{Z}^{\kappa^{2}-4\kappa+5}$ as the stack of all $z_{ij}$ and $\lambda$ in a column vector, the set of constraints can be written as $A_{\text{ineq}}\mathbf{x}\leq\mathbf{b}_{\text{ineq}}$, with proper $A_{\text{ineq}}$ and $\mathbf{b}_{\text{ineq}}$.
@@ V-A2 Truncation-Based Integer Formulation
In this way, the objective function $\zeta(\mathbf{x})$ is evaluating a duplicate-free variable length sequence using the baseline fitness function (5). Although this method lacks the mathematical elegance of the permutation-based integer formulation, it provides a practical and straightforward solution for handling variable-length and unique waypoint assignment sequences without increasing the dimension of the optimization space or adding nonlinear constraints to handle, for instance, the duplicate-free requirement using a penalty method.
@@ V-A3 Benchmarking Results
Both the permutation-based and truncation-based integer formulations are expressed in a MINLP framework, enabling the use of solvers designed for such problems. Since our algorithm was developed in MATLAB, we compare it against solvers available in the same environment—specifically, MATLAB’s Genetic Algorithm and Surrogate Optimization tools, the two methods currently capable of handling MINLP within MATLAB (R2024b). Therefore, we have four methods to compare against our algorithm, namely, the Permutation-based Integer Formulation solved by the MATLAB’s GA (PIFGA), the Permutation-based Integer Formulation solved by the Surrogate Optimization (PIFSO), the Truncation-based Integer Formulation solved by the MATLAB’s GA (TIFGA), and the Truncation-based Integer Formulation solved by the Surrogate Optimization (TIFSO).
@@
Figure 19: Benchmarking results sorted by the number of intermediate waypoints. In the case of 55 waypoints, the PIFGA method is out of the plot as its BestCost exceed 5000, and the PIFSO method does not produce results due to its inability to handle this number of parameters. Very high values of the BestCost indicate that the penalty function is excessively large, indicating that no feasible solution was found.
@@ VI Limitations and Future Work
Another area for improvement is the handling of dynamic obstacles. While the algorithm effectively navigates predefined, static obstacle environments, real-time adaptation to dynamic obstacles requires further development. Integrating predictive models or reactive obstacle avoidance techniques could enhance robustness in dynamic settings. Similarly, real-time constraint handling remains an open challenge. In the current framework, constraints are validated offline during the candidate solution evolution phase, producing a trajectory that maximizes total reward. However, once execution begins, the algorithm does not re-plan the trajectory, leaving the actual control strategy—whether predictive, reactive, or otherwise—independent of the offline planning process. While this flexibility supports a broad range of controllers, real-time corrections to handle unexpected violations warrant further investigation.
@@
To achieve fully autonomous robot operation, with human intervention only when necessary, a promising future direction involves inferring human intentions from demonstrations using data-driven techniques such as inverse optimal control [41] or inverse reinforcement learning [42]. Once the intention is inferred, the robot could autonomously adjust its decision-making to align with the human’s intent [3, 4], even in real-time [42], enabling more responsive and adaptive interactions.
@@ VII Conclusions
To evaluate algorithmic efficiency, we reformulated the problem into two distinct MINLP approaches: one based on permutation and the other on truncation. We integrated MATLAB’s (R2024b) surrogate optimization and genetic algorithm into these formulations and benchmarked them against the proposed algorithm. The results demonstrate that the proposed algorithm outperforms nearly all scenarios and solvers, with the exception of the permutation-based formulation solved by the genetic algorithm for the case of 10 waypoints. Furthermore, the proposed algorithm exhibited superior scalability, consistently achieving better solutions in scenarios with a high number of waypoints. Furthermore, the algorithm demonstrated a time complexity of approximately $\mathcal{O}(n^{2})$.

### v2

@@ IV-B1 Ground Vehicle
The maximum available reward is given by $\sum_{i=2}^{11}w_{\boldsymbol{q}_{i}}=110$. Due to the distance constraint, only a subset of the waypoints can be reached, resulting in a maximum collected reward of $92$. The result is illustrated by the trajectory shown in Fig. 15b. Hypothetically, instead of squeezing itself through two obstacles to reach the waypoint with reward 8, it reaches the waypoints of reward 14, 12, and 18, keeping the other waypoints in the same order. In this case, the total collected reward would increase to 96, but the distance traveled would be 8.5 m $>d_{\mathrm{max}}$.
@@ IV-B2 Quadrotor
The geometric environment and obstacles are shown in Fig. 16. There is no distance constraint but a mission time window $t_{\mathrm{max}}=30$ s. The speed constraint is $|\dot{\boldsymbol{x}}|\leq 0.3$ m/s, and the maximum linear acceleration is $0.5$ m/s2. In total, 11 intermediate waypoints are distributed, with rewards starting at $w_{\boldsymbol{q}_{2}}=2$ and increasing in increments of 2, culminating in $w_{\boldsymbol{q}_{12}}=22$. Only the initial position is specified, while the final position is left unconstrained. The optimal trajectory, shown in Fig. 16, achieves a total reward of $96$ within a time of $29.8$ s.
@@ IV-B3 Quadruped
We demonstrate our approach using the Unitree Go1 quadruped robot. The robot is equipped with two front and two rear legs, each having three joints with specific angular constraints. To the best of the authors’ knowledge, no differentially flat mapping from the states and inputs of the robot’s contact-aware multi-rigid-body dynamic model to its outputs has been presented in the literature. Since the proposed algorithm emphasizes high-level task allocation and planning, we utilize a simplified kinematic model that takes forward-backward speed, left-right speed, and orientation as inputs, following a similar concept in the literature [40]. Although this simplification limits more complex planning, such as navigating uneven terrains [41], the use of hierarchical planning with models of varying fidelity remains effective for task planning [40]. In this high-level framework, we consider the state $\boldsymbol{x}=[x_{1}\ x_{2}\ \theta]^{T}$ in a global reference frame, as indicated in Fig. 17, and we define the flat output as $\boldsymbol{y}=\boldsymbol{x}$. Then, the map from the flat output to the state is the identity, and the map from the flat output and its derivatives to the flat inputs is
@@ V-A1 Permutation-Based Integer Formulation
for $i=2,3,\dots,\kappa-1$ (waypoints), and $j=2,3,\dots,\kappa-1$ (positions in the sequence). We exclude the fixed initial and final waypoints. To have each waypoint assigned to at most one position, one can use $\sum_{j=2}^{\kappa-1}z_{ij}\leq 1,\forall i\in\{2,3,\dots,\kappa-1\}$, and similarly for each position to have at most one waypoint assigned $\sum_{i=2}^{\kappa-1}z_{ij}\leq 1,\forall j\in\{2,3,\dots,\kappa-1\}$. With these two constraints, we assure there are no repetitions, and the robot will reach a single waypoint at a time. Moreover, since we want to force variable length sequences, we include the total number of assignments (intermediate waypoints) as a decision variable $\lambda$, and we constraint the waypoint assignments by $\sum_{i=2}^{\kappa-1}\sum_{j=2}^{\kappa-1}z_{ij}-\lambda\leq 0$. Finally, defining the decision variable $\mathbf{x}\in\mathbb{Z}^{\kappa^{2}-4\kappa+5}$ as the stack of all $z_{ij}$ and $\lambda$ in a column vector, the set of constraints can be written as $A_{\text{ineq}}\mathbf{x}\leq\mathbf{b}_{\text{ineq}}$, with proper $A_{\text{ineq}}$ and $\mathbf{b}_{\text{ineq}}$.
@@ V-A2 Truncation-Based Integer Formulation
In this way, the objective function $\zeta(\mathbf{x})$ evaluates a duplicate-free variable-length sequence using the baseline fitness function (5). Although this method lacks the mathematical elegance of the permutation-based integer formulation, it provides a practical and straightforward solution for handling variable-length and unique waypoint assignment sequences without increasing the dimension of the optimization space or adding nonlinear constraints to handle, for instance, the duplicate-free requirement using a penalty method.
@@ V-A3 Benchmarking Results
Both the permutation-based and truncation-based integer formulations are expressed in a MINLP framework, enabling the use of solvers designed for such problems. Since our algorithm was developed in MATLAB, we compare it against solvers available in the same environment—specifically, MATLAB’s Genetic Algorithm and Surrogate Optimization tools, the two methods currently capable of handling MINLP within MATLAB (R2024b). Therefore, we have four methods to compare against our algorithm, namely, the Permutation-based Integer Formulation solved by MATLAB’s GA (PIFGA), the Permutation-based Integer Formulation solved by the Surrogate Optimization (PIFSO), the Truncation-based Integer Formulation solved by MATLAB’s GA (TIFGA), and the Truncation-based Integer Formulation solved by the Surrogate Optimization (TIFSO).
@@
Figure 20: Benchmarking results sorted by the number of intermediate waypoints. In the case of 55 waypoints, the PIFGA method is out of the plot as its BestCost exceeds 5000, and the PIFSO method does not produce results due to its inability to handle this number of parameters. Very high values of the BestCost indicate that the penalty function is excessively large, indicating that no feasible solution was found.
@@ VI Limitations and Future Work
Another area for improvement is the handling of dynamic obstacles. While the algorithm effectively navigates predefined, static obstacle environments, real-time adaptation to dynamic obstacles requires further development. Integrating predictive models or reactive obstacle avoidance techniques could enhance robustness in dynamic settings. As a future improvement, we propose that when the environment changes during mission execution, seeding the initial population with the offline-computed reference solution can enable the algorithm to rapidly converge to a new, feasible, albeit suboptimal, solution. Similarly, real-time constraint handling remains an open challenge. In the current framework, constraints are validated offline during the candidate solution evolution phase, producing a trajectory that maximizes total reward. However, once execution begins, the algorithm does not re-plan the trajectory, leaving the actual control strategy—whether predictive, reactive, or otherwise—independent of the offline planning process. While this flexibility supports a broad range of controllers, real-time corrections to handle unexpected violations warrant further investigation.
@@
To achieve fully autonomous robot operation, with human intervention only when necessary, a promising future direction involves inferring human intentions from demonstrations using data-driven techniques such as inverse optimal control [43] or inverse reinforcement learning [44]. Once the intention is inferred, the robot could autonomously adjust its decision-making to align with the human’s intent [3, 4], even in real time [44], enabling more responsive and adaptive interactions.
@@ VII Conclusions
To evaluate algorithmic efficiency, we reformulated the problem into two distinct MINLP approaches: one based on permutation and the other on truncation. We integrated MATLAB’s (R2024b) surrogate optimization and genetic algorithm into these formulations and benchmarked them against the proposed algorithm. The results demonstrate that the proposed algorithm outperforms nearly all scenarios and solvers, except the permutation-based formulation solved by the genetic algorithm for the case of 10 waypoints. Furthermore, the proposed algorithm exhibited superior scalability, consistently achieving better solutions in scenarios with a high number of waypoints. Furthermore, the algorithm demonstrated a time complexity of approximately $\mathcal{O}(n^{2})$.